OpenStudy (anonymous):

Given cos 2θ=-1/2 and 2θ ends in quadrant II. Find cotθ

OpenStudy (anonymous):

1) solve for \[\theta\] \[2\Theta=4\pi/3\] \[\theta=2\pi/3\] it's means that: \[\cot \theta =\cot (2\pi/3)=(-1/2) /(\sqrt{3}/2) = -1/\sqrt{3}\]